OpenStudy (anonymous):

Algebra II Question. Please help ASAP. I need x,y and z. Thank you in advance! (This is elimination) 4x+y+3z=0 2x-2y-z=10 3x-2y+2z=11

OpenStudy (campbell_st):

well the most obvious is to eliminate y so equation 1 times 2 + equation 2 8x + 2y + 6z = 0 2x - 2y - z = 10 -------------- 10x + 5z = 10 divide each term by 5 and you get 2x + z = 2 call it equation (4) then 2 times equation 1 + equation 3 8x + 2y + 6z = 0 3x - 2y + 2z = 11 ----------------- 11x + 8z = 11 equation (5) so you are now solving 2 equations in 2 unknowns 2x + z = 2 equation 4 11x + 8z = 11 equation 5 so equation 1 times 8 and subtract equation 5 will eliminate z 16x + 8z = 16 - 11x + 8z = 10 ---------------- 5x = 5 so x = 1 you can now go and use substitution to find y and z. hope it helps

OpenStudy (anonymous):

I don't know how to find y and z. This is so confusing for me

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well take equation 4 and substitute x =1 to find the value of z when you get z go to any of the initial equations and subsitute x = 1, z = your answer to find y

OpenStudy (anonymous):

would z =0? @campbell_st

OpenStudy (campbell_st):

thats correct now y

OpenStudy (anonymous):

Where am I supposed to plug in z? @campbell_st

OpenStudy (campbell_st):

use any of the initial equations you are given, substitute x = 1, z = 0 to find y. choose which ever you feel comfortable with.

OpenStudy (anonymous):

y=-4? @campbell_st

OpenStudy (campbell_st):

yes it is well done

OpenStudy (anonymous):

I put that in and it told me I was wrong? @campbell_st

OpenStudy (anonymous):

Nevermind I got it! I wrote it wrong! Thank you for your help! :) @campbell_st